OpenStudy (anonymous):

How do you do Elimination Using Multiplication? An example problem could be like: 2x+5y=3 -x+3y=-7 can you walk me through please?

OpenStudy (luigi0210):

do you mean like solve for x and y?

hero (hero):

1. Multiply the second equation by 2 2. Add both equations together.

OpenStudy (anonymous):

But how do you figure out what to multiply it by?

OpenStudy (luigi0210):

find a number that cancels out the other

OpenStudy (luigi0210):

leaving just one variable

OpenStudy (anonymous):

But how?

hero (hero):

You figure it out by observing which variable is easiest to eliminate. Then you use the rules of algebra to make the co-efficients of the variable the same for both equations. Afterwards, you subtract one equation from the other.

OpenStudy (anonymous):

I get everything except im still struggling to understand the concept of how you find the number that cancels it out can u quick tell me step by step how to solve this equation: 7x+15y=32 1x-3y=20

hero (hero):

Thought process: 1. The coefficients of variable x for both equations are 2 and -1. 2. The coefficients of variable y are 5 and 3. 3. It is easier to eliminate variable x because -1 x 2 = -2, the opposite of 2. 4. Which means if I multiply the second equation by 2, I would be able to eliminate the x variable by adding both equations together.

hero (hero):

Do the same thought process for the second problem you are working on.

OpenStudy (anonymous):

Could you still take me through the steps on the one i just posted? Sorry...

OpenStudy (anonymous):

Nvm

hero (hero):

7x+15y=32 1x-3y=20 1. Ask yourself, "How can I make the coefficents of x the same for both equations?". The answer is to multiply the second equation by 7. Afterwards, to eliminate x, subtract one equation from the other.

OpenStudy (anonymous):

Hey i actually did tht one correctly..

hero (hero):

Alternatively, you could ask yourself: "How can I make the coefficents of y the same?". The answer is to multiply the second equation by 5, then add both equations together.

OpenStudy (anonymous):

Say whattt

OpenStudy (luigi0210):

multiply then add them together

hero (hero):

7x+15y=32 1x-3y=20 Solution 1: Multiply the second equation by 7: 7x + 15y = 32 7x - 21y = 140 Subtract the second equation from the first: 36y = -108 y = -108/36 y = -3 Solution 2: Multiply the second equation by 5: 7x+15y=32 5x- 15y=100 Add both equations together: 12x = 132 x = 132/12 x = 11

OpenStudy (anonymous):

So y is -3 and x is 11?

OpenStudy (anonymous):

@mintymint Do you actually get how to solve this problems now? Or are you still stuck about what numbers to multiply by and so on?

OpenStudy (anonymous):

I think i sorta get it. So do you find the least common denominator and use the numbers you you would multiply by to get the number as the coeffients? Just say yes or no.

hero (hero):

LCD only applies to fractions.

OpenStudy (anonymous):

Whats lcd?

hero (hero):

Least Common Denominator

OpenStudy (anonymous):

This is confusing.

OpenStudy (luigi0210):

all you do is take both problems and solve them for one variable, then plug in that number for the variable to get the remaining variable

hero (hero):

It's so obvious, it's confusing for you @mintymint

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Thank you.

OpenStudy (luigi0210):

Try it out, if your still confused I could try and help

hero (hero):

@mintymint, please check your messages.

OpenStudy (texaschic101):

wouldn't you multiply the second equation by a - 7 instead of 7 ?

hero (hero):

@texaschic101, my steps speak for itself.

OpenStudy (anonymous):

Why would you multiply with negatives, when you can have a higher chance of making mistakes?

OpenStudy (texaschic101):

but your steps won't cancel out the x

OpenStudy (anonymous):

YOu could subtract the two equations?

hero (hero):

Read the steps again @texaschic101

OpenStudy (texaschic101):

oh, I didnt subtract, I added......can it be done either way ?

OpenStudy (anonymous):

Yes, but if you make more mistakes when multiplying than subtracting, then don't come back and tell me "you're right."

hero (hero):

It's the same process.

hero (hero):

There are a few reasons why I explain it that way: 1. I am not a big fan of multiplying (or dividing ) by negatives. 2. I consider my method a way of explaining every step more explicitly. 3. I feel it is clearer to understand the process

OpenStudy (texaschic101):

I dont understand how you got 36y in your first solution

OpenStudy (texaschic101):

oh wait, I think I know......-21y - (-15y) never mind I see

OpenStudy (texaschic101):

oh wait....15y - (-21y)...sorry I understand now

hero (hero):

Well, that's good.

OpenStudy (texaschic101):

sorry about the trouble.....you are a patient person

hero (hero):

Patience suggests that I was waiting for something, but I wasn't. The understanding of it was totally up to you.

OpenStudy (texaschic101):

smart too

OpenStudy (anonymous):

I think i get it now. Just to be clear, when i grow up, i am not going to be an accountant or doctor. Haha.

hero (hero):

You cannot avoid math. Math is used to create everything.

OpenStudy (anonymous):

Yeah but if your in like the fashion industry, sure there is math, but it doesnt involve all this crap.

hero (hero):

I met Math. She is a nice person. She just wants you to get to know her. She gets sad every time someone says "I hate Math" or every time someone tries to avoid her. She's a really cool person.

OpenStudy (anonymous):

But in all seriousness, yeah math is important, but lets face it. Half of the stuff we learn were not even going to use. It doesnt mean i dont care, it just means its true. Were just two opposite people thats all.

OpenStudy (anonymous):

P.s. thanks for the storybut im not 5 lol

hero (hero):

Writing involves the ability to use logic and describe events in sequence, as well as critical thinking. Funny thing is, math requires those same exact things. Writing and Math is not as far off as you think. In fact, many English sentences can be directly translated to mathematical equations.

OpenStudy (anonymous):

Well yes, or you can just tell a story.

hero (hero):

But Math also tells a story...just in a different language. In fact, math is considered a universal language. Other languages, are limited.

OpenStudy (anonymous):

But it cant solve everything.

hero (hero):

Math is about the closest thing there is to solving everything.

hero (hero):

Math is solely responsible for the advancement of humanity.

hero (hero):

Without math, we're still living in the stone ages.

OpenStudy (anonymous):

I wonder how math begun. Isnt it amazing how ipads are built from technically natuaral resources?

hero (hero):

If you had two worlds: One knew math, the other knew writing....which one would advance the fastest?

OpenStudy (anonymous):

Well in order to figure out.math you must communicate

hero (hero):

Math begun with the need to keep track of things. Like early humans wanted to make sure that when they gathered things, they could make sure that still had the same amount of stuff tomorrow that they had yesterday.

hero (hero):

Math IS a form of communication.

hero (hero):

It is a language within itself. It uses symbols, visualization, logic, sequence, structure, and form to express itself.

hero (hero):

...and sentences.

OpenStudy (anonymous):

But yet a not even a math equation can solve how this universe begun. If not a universe nothing would exist, not even colors

OpenStudy (anonymous):

But bycommunicating in words is how onecomes to understand a concept of math

hero (hero):

Actually, math is the best tool we have for figuring out how the universe begun. Tools are being used as we speak to figure it out, and I must admit, they're closer than ever to figuring it out. I just hope they don't destroy the planet trying to.

hero (hero):

You don't always have to communicate in words to communicate. Remember, there are several forms of communication. You can communicate things to someone without saying a word to them.

OpenStudy (anonymous):

Then explain to me why we have languages. Why is there not a cure for disease if we are so far advanced. Is there some mathimatical concept we missed?

OpenStudy (anonymous):

By disease i mean cancer. I mean my mom a few years back had cancer so i.get it.

hero (hero):

There is a cure for disease, and believe it or not, it is mathematical in nature.

OpenStudy (anonymous):

What about cancer, what makes it so different

hero (hero):

First of all, you have to consider the fact that we ourselves are biological computers with automated biological processes. There are two such automated biological processes that occur in humans of which we are duly familiar: heart beat and thought. Of course these are only two things most familiar to every one, the operations of which are similar in nature to electro-mechanical devices such as generators and water pumps. As far as cancer is concerned, it is simply the malfunction of cells. Cells need nutrients just like we do and when we don't provide them with nutrients and proper cleansing, they die. Those are the two things required for cellular function: 1. Nutrients 2. Cleansing The reason why people get cancer (or die) is because they neglect one or both of these things.

hero (hero):

For nutrients, it has been discovered that the main source of cellular nutrition are what is known as "cellular salts". These cellular salts provides the basic nutrients needed for healthy cellular metabolism. For cleansing, it basically comes down to oxygen. Cells need oxygen to "breathe" just like we do. When blockages occur in the body due to toxic build-up, these toxins can block the flow of oxygen to certain parts of the body or certain organs which can over time result in the cellular malfunction we call cancer. Cancer manifests itself in many different forms, but the cause is essentially the same.

hero (hero):

The entire body both micro and macro processes consists of mostly automated processes controlled by the brain. This is why we can consider ourselves biological computers. Our body conducts seemingly automated processes with nutrients such as foods, vitamins, and minerals being the source of energy.

OpenStudy (anonymous):

Or, maybe all parts of our body are branches of living things. We are just the main control.

hero (hero):

We are the whole organism, a collection of biological entities working synergistically and dependently to help the whole work. I think you should research cellular salts. You could learn a lot.

OpenStudy (texaschic101):

wow.......this is deep !!!

OpenStudy (anonymous):

Its also hard for cancer to be cured i suppose because all of our genetics are made up uniqly. This may explain why people that have smoked sometimes dont get cancer or have not as opposed to those who dont but still get it.

hero (hero):

Cell Salt deficiency is one of the main causes of most diseases in humans. In many cases, call salt deficiency creates an environment that allows disease to proliferate.

hero (hero):

Cellular Salts are important because they both provide cells with nutrients and help to cleanse the body at the cellular level. There are 12 different cellular salts that correspond with 12 different regions of the body and each are responsible for cleansing and providing nutrients to the cells in their respective regions.

OpenStudy (anonymous):

It kind of reminds me of if you leave perishable foods out for too long, they attract mold. Kind of the same concept, except a bit different.

hero (hero):

I leave perishable foods out all the time, particularly fruits and vegetables. However, the difference is I wrap them with suran wrap to protect them from oxidation.

hero (hero):

Very tightly

OpenStudy (anonymous):

Ps how old are you you could be a doctor

hero (hero):

I don't go around openly professing my age.

OpenStudy (texaschic101):

doctor or biologist ...told you hero was smart

hero (hero):

As far as I'm concerned, I'm ageless

OpenStudy (anonymous):

So im just takin a shot in the darke but maybe the reason youre also likely to get sick or something if you constantly eat bad things is because if its natural like fruits those are easy for your body to process but if its artificial its takes longer for your body to identify it so the air you intake has an effect on it. Probably wrong but hey atleast im creative!

OpenStudy (anonymous):

And i assume youre an adultof some sort maybe? Well im 14.

hero (hero):

Also, another thing about the suggestion that humans are biological computers. When programmers program computers, they write instructions in the form of code for the computer program to execute. Then the computer executes or interprets and performs whatever it is instructed to do. When we as humans eat food, the food has its own form of instructions or code embedded within it. Of course the code is vitamins, nutrients, minerals, acids, salts, etc. The brain takes all of the "code" from the food and disseminates (interpreting) it to the appropriate areas of the body which in turn processes it for its respective purpose, which is essentially providing nutrients for proper function. However, just like a computer can contract a virus, we as humans can also contract a virus. A virus is like "bad code" that can destroy a computer program. Bad food or toxins is like bad code. It contains things in it that can destroy us. And just like there are many kinds of computer viruses, there are many kinds of viruses the human body can contract in many different forms and of course not just limited to food.

OpenStudy (anonymous):

Were part of the matrix. Ha, no.

OpenStudy (anonymous):

We can also sleep and turn on like compiters and need to "recharge" like computers.